### v1

# Dynamic Scene Understanding through Object-Centric Voxelization and Neural Rendering

Abstract
Learning object-centric representations from unsupervised videos is challenging. Unlike most previous approaches that focus on decomposing 2D images, we present a 3D generative model named DynaVol-S for dynamic scenes that enables object-centric learning within a differentiable volume rendering framework. The key idea is to perform object-centric voxelization to capture the 3D nature of the scene, which infers per-object occupancy probabilities at individual spatial locations. These voxel features evolve through a canonical-space deformation function and are optimized in an inverse rendering pipeline with a compositional NeRF. Additionally, our approach integrates 2D semantic features to create 3D semantic grids, representing the scene through multiple disentangled voxel grids. DynaVol-S significantly outperforms existing models in both novel view synthesis and unsupervised decomposition tasks for dynamic scenes. By jointly considering geometric structures and semantic features, it effectively addresses challenging real-world scenarios involving complex object interactions. Furthermore, once trained, the explicitly meaningful voxel features enable additional capabilities that 2D scene decomposition methods cannot achieve, such as novel scene generation through editing geometric shapes or manipulating the motion trajectories of objects.
1 Introduction
Unsupervised learning of the physical world is of great importance but challenging due to the intricate entanglement between the spatial and temporal information [1, 2, 3]. Existing approaches primarily leverage the consistency of the dynamic information across consecutive video frames but tend to ignore the 3D nature, resulting in a multi-view mismatch of 2D object segmentation [4, 5, 6]. This paper explores a novel research problem: unsupervised 3D dynamic scene decomposition. Unlike previous 2D approaches, our method naturally ensures 3D-consistent scene understanding and achieves an explicit knowledge of the object geometries and physical interactions in the dynamic scenes.
We build the model, DynaVol-S, in an inverse graphics framework based on neural radiance field (NeRF) to effectively leverage the multi-view video frames collected from a monocular camera. However, the classic canonical-space rendering pipeline does not support object-centric learning. We identify two key challenges: First, how to decouple the underlying dynamics of each object from its visual appearance and capture the local geometric structures of each object. Second, how to integrate the global semantic priors of the scene, typically extracted from 2D foundation models, into 3D volume rendering, which is crucial for inferring 3D geometries given partial observations, complex real-world structures, and severe occlusions.
To address the first challenge and integrate object-centric learning into the volume rendering method, we propose a new form of object-centric representation that captures the local spatial structures of each object using time-aware voxel grids. This object-centric voxelization method explicitly allows us to infer a distribution of occupancy probabilities over objects at individual spatial locations, thereby naturally facilitating 3D-consistent scene decomposition. Unlike existing neural rendering techniques for dynamic scenes, such as TineuVox [7], our approach includes two key modifications in the model design (see Fig. 1). First, it exploits the canonical-space deformation network to transform the voxelized object-centric representations over time. Second, based on the transformed occupancy probabilities per object, it employs a compositional NeRF for view-dependent image rendering, which also relies on a set of object-centric latent codes to represent the time-invariant features of each object. Notably, in contrast to prior research that focuses on 3D dynamic scene reconstruction [7, 8, 9], our unsupervised voxelization approach provides two additional advantages beyond novel view synthesis. First, it allows for fine-grained separation of object-centric information in 3D space, enabling a better understanding of the scene. Second, it enables direct scene editing (e.g., object removal, replacement, and trajectory modification) that is not feasible in existing video decomposition methods. This is done by directly manipulating the voxel grids or the learned deformation function without the need for further training.
For the second challenge of semantics integration, we observe that extending our previous work of DynaVol [10] to real-world data poses significant challenges due to the increased complexity of color and geometry patterns and the limited range of available view directions. To address these issues, we present a new method that leverages 2D semantic feature maps extracted from an image encoding network trained on large-scale datasets. We develop a novel semantic volume slot attention module to re-project the 2D feature maps onto 3D voxel grids. This design enables our model to correlate the object-centric voxel grids with the pre-learned 2D feature maps, incorporating common knowledge of object geometries. This is particularly beneficial for handling previously unseen object structures, enhancing the model’s ability to interpret complex real-world scenes.
Figure 1: DynaVol-S explores unsupervised object-centric decomposition in 3D dynamic scenes within an inverse rendering framework. Unlike previous canonical-space neural rendering techniques, such as TineuVox [7], our approach integrates voxelized object-centric representations, which achieves an explicit understanding of the object geometries and physical interactions in the dynamic scenes and further facilitates downstream tasks such as scene editing.
In our experiments, we initially evaluate DynaVol-S on simulated 3D dynamic scenes that contain different numbers of objects, diverse motions, shapes (such as cubes, spheres, and real-world shapes), and materials (such as rubber and metal). In the simulated dataset, we directly assess the performance of DynaVol-S for scene decomposition by projecting the object-centric volumetric representations onto 2D planes and comparing it with existing approaches, such as SAVi [11], uORF [12], and MovingParts [13]. Subsequently, we conduct a series of experiments on real-world scenes. Specifically, we compare DynaVol-S against NeuralDiff [14], HyperNeRF [8], TineuVox [7], and $\text{D}^{2}$eRF [9] in novel view synthesis tasks, and against OCLR [15] and ClipSeg [16] in scene decomposition tasks. These experiments validate the effectiveness of DynaVol-S, highlighting the necessity of incorporating semantic features.
In summary, DynaVol-S provides an early exploration of using NeRF-based models to tackle the research problem of unsupervised object-centric decomposition of 3D dynamic scenes. We extend our previous studies presented at ICLR 2024 in the following aspects:
  • Setup: We adopt a more affordable experimental setting, requiring only training data captured by a monocular camera instead of using multiple viewpoints at the initial timestamp.
  • Method: We introduce semantic features to enhance object-centric representation learning, significantly improving performance in real-world scenes.
  • Experiments: We incorporate more real-world scenes and compare with more recent approaches, including TineuVox, MovingParts, OCLR, and ClipSeg, to demonstrate the superiority of DynaVol-S in more challenging scenarios.
2 Problem Setup
We assume a set of RGB images $\{\mathbf{I}_{t},\mathbf{T}_{t}\}_{t=1}^{T}$ collected with a moving monocular camera, where $\mathbf{I}_{t}\in\mathbb{R}^{H\times W\times 3}$ are images acquired under camera poses $\mathbf{T}_{t}\in\mathbb{R}^{4\times 4}$, and $T$ is the length of the sequential video frames. The goal is to decompose each object in the scene by harnessing the underlying space-time structures and semantics present in the visual data without additional supervision.
Please note that, unlike the previous work of DynaVol presented at ICLR 2024, we no longer use multiple viewpoints at the initial timestamp. We consider scenarios where (i) the training images are from viewpoints randomly sampled from the upper hemisphere, following the setup in D-NeRF [17], and (ii) a monocular camera captures the training images along a smooth moving trajectory, which is easily achievable in real-world applications.
3 Method
In this section, we first review the original neural radiance field (NeRF) methodologies (Sec. 3.1). Next, we introduce the concept of object-centric voxel representations and discuss the overall framework of our approach (Sec. 3.2). Subsequently, we present the details of each network component in the proposed DynaVol and DynaVol-S (Sec. 3.3). Finally, we present the three-stage training procedure of our approach (Sec. 3.4).
3.1 Preliminaries on NeRF
The Neural radiance field models 3D scenes by mapping the 3D coordinates $\mathbf{x}$ and view directions $\mathbf{d}$ to the corresponding density $\sigma$ and color $\mathbf{c}$ using a neural network $\Phi$:
$\begin{split}(\sigma,\mathbf{c})=\Phi(\mathbf{x},\mathbf{d}).\end{split}$ (1)
The outputs $\sigma$ and $\mathbf{c}$ are then transformed to the pixel values on a 2D image using volume rendering techniques [18]. Specifically, given a ray $\mathbf{r}=\mathbf{o}+t\mathbf{d}$ emitted from $\mathbf{x}$ to a specific pixel in the image, where $\mathbf{o}$ and $\mathbf{d}$ are respectively the origin and direction of the ray and $t$ is the distance from $\mathbf{x}$ to the camera. The color of the pixel is determined by the integral of the $(\sigma_{i},\mathbf{c}_{i})$ values of all the points along the ray within a predefined near bound $t_{n}$ and a far bound $t_{f}$:
$\begin{split}C(\mathbf{r})=&\sum_{i=1}^{P}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right)\mathbf{c}_{i},\\ T_{i}=&\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right),\end{split}$ (2)
where $P$ denotes the number of sampled points, and $\delta_{i}$ denotes the distance between the $i$-th and $(i+1)$-th points. In this work, we learn object-centric representations that evolve over time to drive a compositional NeRF renderer.
Figure 2: DynaVol-S consists of three groups of network components: the bi-directional dynamics modules ($f_{\psi}$, $f_{\xi}^{\prime}$), the volume slot attention module, and neural rendering modules based on 3D and 4D voxels respectively. The training scheme involves a warmup stage (top), an object-centric voxel grids initialization stage (middle, whose pseudocode is given in Alg. 1), and a multi-grids joint optimization stage (bottom). For better clarity, we present the symbols and their corresponding descriptions in Table I.
TABLE I: Notations.
Symbol Description Symbol Description
$\mathcal{V}^{\text{M}}$ Voxel grids for occupancy probabilities per object $\widehat{C}$ Rendered 2D pixel colors
$\mathcal{V}^{\text{D}}$ Voxel grids for opacity $C$ Observed 2D pixel colors
$\mathcal{V}^{\text{C}}$ Voxel grids for color-related features $\mathcal{A}$ Semantic probabilities
$\mathcal{V}^{\text{S}}$ Voxel grids for semantic features $\widehat{\mathcal{A}}$ Rendered 2D semantic probabilities
$\sigma$ Opacity $\widehat{F}$ Rendered 2D semantic features
$\mathbf{c}$ Predicted color $F$ 2D semantic features extracted by DINOv2
$\mathbf{f}^{c}$ Color-related features $\{p_{n}\}_{n=1}^{N}$ Occupancy probabilities per object
$\mathbf{f}^{s}$ Semantic features $\widehat{P}$ Rendered 2D occupancy probabilities
$N$ Number of latent codes $\bar{P}$ Pseudo-labels obtained by CRF
$\{\mathbf{z}_{n}\}_{n=1}^{N}$ Object-centric latent codes $f_{\psi},f_{\xi}^{\prime}$ Bi-directional deformation networks
$\mathbf{S}$ / $\{\mathbf{s}_{n}\}_{n=1}^{N}$ Object-centric slot features $N_{\phi},N_{\phi^{\prime}}$ Compositional/Non-compositional NeRF
3.2 Object-Centric Voxelization
DynaVol-S is trained in an inverse graphics framework to synthesize $\{\mathbf{I}_{t}\}_{t=1}^{T}$ without any further supervision. Formally, the goal is to learn an object-centric projection of $(\mathbf{x},\mathbf{d},t,\{\mathbf{z}_{n}\}_{n=1}^{N},\mathcal{V}^{\text{M}},\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}})\xrightarrow[]{}(\mathbf{c},\sigma)$. $\mathbf{x}=(x,y,z)$ is a 3D point, $\{\mathbf{z}_{n}\}_{n=1}^{N}\in\mathbb{R}^{N\times D_{z}}$ denotes a set of learnable latent codes associated with individual objects, and $N$ is the predefined number of the latent codes that is assumed to be larger than the number of objects. $\mathcal{V}^{\text{M}},\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}}$ are voxelized features, which respectively represent occupancy probabilities per object, opacity, and color-related features.
To understand the scene from an object-centric perspective, we extend the classic approach of using 3D voxel grids for volume rendering by incorporating 4D voxel grids, represented as $\mathcal{V}^{\text{M}}\in\mathbb{R}^{N\times N_{x}\times N_{y}\times N_{z}}$, where $(N_{x},N_{y},N_{z})$ denote the spatial resolutions. The newly added dimension in the expanded voxel features capture the occupancy probabilities for each object at the grid vertices. Therefore, the occupancy probabilities $\{p_{n}\}_{n=1}^{N}$ at any given 3D point within the grid cells are efficiently retrieved using trilinear interpolation:
$\mathrm{Interp}(\mathbf{x},\mathcal{V}^{\text{M}}):\ (\mathbb{R}^{3},\mathbb{R}^{N\times N_{x}\times N_{y}\times N_{z}})\xrightarrow{}\mathbb{R}^{N}.$ (3)
Following this, we apply the Softmax activation function to the interpolated values to normalize $\{p_{n}\}_{n=1}^{N}$, ensuring they sum to one and are proportionately scaled.
In addition to $\mathcal{V}^{\text{M}}$, our approach employs three other voxelized representations that are independent of the object index: $\mathcal{V}^{\text{D}}\in\mathbb{R}^{1\times N_{x}\times N_{y}\times N_{z}}$ for modeling the opacity $\sigma$, $\mathcal{V}^{\text{C}}\in\mathbb{R}^{D_{c}\times N_{x}\times N_{y}\times N_{z}}$ for modeling color-related features $\mathbf{f}^{c}$, and $\mathcal{V}^{\text{S}}\in\mathbb{R}^{D_{s}\times N_{x}\times N_{y}\times N_{z}}$ for modeling semantic features of the scene. The incorporation of semantic features is a pivotal contribution of the improved DynaVol-S. We optimize the initial states of $\{\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}},\mathcal{V}^{\text{S}},\mathcal{V}^{\text{M}}\}$ and learn a deformation module $f_{\psi}(\cdot)$ to transform them over time in the canonical space.
Apart from the voxelized representations that present the local geometry and appearance of the scene, our approach also learns a set of object-centric latent codes $\{\mathbf{z}_{n}\}_{n=1}^{N}$ to capture the global object-specific information and particularly use them in the NeRF renderer. We set $\{\mathbf{z}_{n}\}_{n=1}^{N}$ as time-invariant features.
3.3 Model Components
As shown in Fig. 2, DynaVol-S consists of three modules:
  1. (i)
    The bi-directional deformation networks, $f_{\psi}$ and $f_{\xi}^{\prime}$, which learn the canonical-space transitions over time.
  2. (ii)
    The semantic volume slot attention module, which learn the semantic features $\mathcal{V}^{\text{S}}$ and disentangle them into object-centric slot features $\mathbf{S}$ respectively. We extract a set of 2D semantic-level features $\{\mathbf{F}_{t}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}\}_{t=1}^{T}$ from $\{\mathbf{I}_{t}\}_{t=1}^{T}$ using a pre-trained 2D feature extractor, e.g., DINOv2 [19] as additional learning objectives for this module.
  3. (iii)
    The neural renders $N_{\phi}$ and $N_{\phi^{\prime}}$. $N_{\phi}$ is a compositional NeRF that jointly uses $\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}},\mathcal{V}^{\text{M}}$, and $\{\mathbf{z}_{n}\}_{n=1}^{N}$ for rendering of each object. $N_{\phi^{\prime}}$ is a non-compositional NeRF only used to initialize the voxel representations.
3.3.1 Bi-Directional Deformation Networks
Canonical-space dynamics modeling. We leverage the canonical space for dynamics modeling, where the canonical coordinates are used in the Hamiltonian formulation of classical mechanics to describe a physical system at any given point in time. Specifically, as illustrated in Fig. 2, we use a dynamics module $f_{\psi}(\mathbf{x}_{i},t)$ to learn the deformation field, where $\mathbf{x}_{i}$ and $t$ are first projected into higher dimensions through positional encoding. This module transforms a 3D point $\mathbf{x}_{i}$ at an arbitrary time to its corresponding position at the initial time by predicting a position movement $\Delta\mathbf{x}_{i}$ and calculating $\mathbf{x}_{i}^{\prime}=\mathbf{x}_{i}+\Delta\mathbf{x}_{i}$. Using canonical-space mechanics, we determine the variations in voxel representations over time. For any given coordinate $(\mathbf{x}_{i},t)$, we have:
$\small\begin{split}\sigma_{i}=\mathrm{Interp}&\left(\mathbf{x}_{i}^{\prime},\mathcal{V}^{\text{D}}\right),\\ \mathbf{f}_{i}^{c}=\mathrm{Interp}\left(\mathbf{x}_{i}^{\prime},\mathcal{V}^{\text{C}}\right)&\oplus\ldots\oplus\mathrm{Interp}\left(\mathbf{x}_{i}^{\prime},\mathcal{V}^{\text{C}}\left[::s_{M}\right]\right).\end{split}$ (4)
The opacity is retrieved through trilinear interpolation. For the color-related voxel features, we adopt a multi-distance interpolation strategy [7], where $\oplus$ denotes concatenating features across various resolutions and $[::s_{M}]$ indicates the stride used to sample the grid at various levels. We set $s_{M}=\{2,4\}$ in our experiments. The retrieved voxel features are then used in the neural renderer.
Forward dynamics modeling. Additionally, we use another dynamics module $f_{\xi}^{\prime}(\mathbf{x}_{i},t)$ to model the forward movement $\Delta x_{i}^{\prime}$ from the initial time to timestamp $t$. This module enables the calculation of a cycle-consistency loss, enhancing the coherence of the learned canonical-space transitions. It provides useful information representing the forward dynamics for the subsequent connect components algorithm in the warmup stage (described in detail in Sec. 3.4), which can improve the initialization of the aforementioned voxel grids.
3.3.2 Semantic Volume Slot Attention
To extract semantic features and leverage them to improve scene decomposition by correlating them with voxel grids, we design a simple yet effective module named semantic volume slot attention.
Semantic feature extractor. We use DINOv2 [19] as the semantic feature extractor and further employ FeatUp [20] to generate higher-resolution feature maps. Notably, both methods are trained in an unsupervised manner. We particularly use the semantic features in real-world scenes, while our previous work on DynaVol does not exploit any semantic features. For synthetic scenes, instead, we observe that our approach can effectively handle the object geometries even in the absence of semantic features.
Volume slot attention. During the warmup phase, we employ the volume slot attention to progressively derive a set of global time-invariant object-level representations from the local, time-varying voxel features. Specifically, we introduce the semantic voxel grids, $\mathcal{V}^{\text{S}}\in\mathbb{R}^{D_{s}\times N_{x}\times N_{y}\times N_{z}}$, and a set object-centric “slots features”, $\mathbf{S}\in\mathbb{R}^{N\times D_{s}}$, borrowed from the prior literature on 2D static scene decomposition [21]. Given a 3D point $\mathbf{x}$ at timestamp $t$, we have:
$\begin{split}\mathbf{f}_{i}^{s}\in\mathbb{R}^{D_{s}}&=\text{Interp}\left(\mathbf{x}_{i}+f_{\psi}(\mathbf{x}_{i},t),\mathcal{V}^{\text{S}}\right)\\ \mathcal{A}_{i}\in\mathbb{R}^{N}&=\mathrm{softmax}_{N}\left(\frac{\mathbf{f}_{i}^{s}\cdot\mathbf{S}^{T}}{\tau}\right)\\ \mathbf{f}_{i}^{s}\in\mathbb{R}^{D_{s}}&=\mathcal{A}_{i}\cdot\mathbf{S},\end{split}$ (5)
where $T$ denotes the transpose of a matrix, $\tau$ is a learnable temperature factor, and $\mathcal{A}_{i}$ represents the semantic probabilities that are further used to initialize the 4D voxel grids, $\mathcal{V}^{\text{M}}$. Unlike prior slot attention methodologies, we do not use query, key, and value projections. This approach improves efficiency and ensures that linear projections do not affect the correlation between features and slots. Subsequently, $\mathbf{f}^{s}_{i}$ and $\mathcal{A}_{i}$ are used to obtain the 2D semantic features and corresponding 2D maps of semantic probabilities with the quadrature rule [18]:
$\begin{split}\widehat{F}(\mathbf{r})&=\sum_{i=1}^{P}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right)\mathbf{f}_{i}^{s}\\ \widehat{\mathcal{A}}(\mathbf{r})&=\sum_{i=1}^{P}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right)\mathcal{A}_{i}.\end{split}$ (6)
We then use a linear layer to map $\widehat{F}(\mathbf{r})$ from $\mathbb{R}^{D_{s}}$ to $\mathbb{R}^{C}$, the channel dimension of the 2D semantic feature maps provided by DINOv2, and compute the reconstruction loss as described in Eq. (10). Additionally, we supervise the accumulated $\widehat{\mathcal{A}}(\mathbf{r})$ using the entropy loss, as detailed in Eq. (10), to further reduce the ambiguities of the learned semantic representations. For an empirical analysis of such model designs, please refer to Sec. 4.5.
3.3.3 Object-Centric Renderer
Previous compositional NeRF models, like uORF [12], typically use MLPs to learn a continuous mapping from sampling point, viewing direction, and slot features to the emitted densities and colors of different objects. In contrast, our neural renderer takes as inputs $\{\mathbf{z}_{n}\}_{n=1}^{N}$, which is a set of learnable parameters representing the global object-centric features. In DynaVol-S, we learn the linear combinations of $\{\mathbf{z}_{n}\}_{n=1}^{N}$ to construct the inputs of the rendering network, such that:
$\begin{split}p_{i}=&\ \mathrm{softmax}(\mathrm{Interp}(\mathbf{x}_{i}+f_{\psi}(\mathbf{x}_{i},t),\mathcal{V}^{\text{M}});\mathrm{Dim}=N),\\ \overline{\mathbf{z}}_{i}=&\sum_{n=1}^{N}p_{i}[n]\cdot\mathbf{z}_{n}.\end{split}$ (7)
Subsequently, we use an MLP to learn object-centric projections $N_{\phi}:(\mathbf{f}_{i}^{c},\mathbf{d},\overline{\mathbf{z}}_{i})\xrightarrow{}\mathbf{c}_{i}$, where $\mathbf{f}_{i}^{c}$ are color-related features from Eq. (4). We apply positional encodings to $\mathbf{f}_{i}^{c}$ and $\mathbf{d}$ before feeding them to the MLP. In the rendering process, we directly query the opacity $\sigma$ from the voxel grids $\mathcal{V}^{\text{D}}$ within the canonical space. We follow Eq. (2) to estimate the pixel color of a sampling ray using the predicted $\mathbf{c}_{i}$ and the retrieved $\sigma_{i}$.
3.4 Training Procedure
DynaVol-S involves three stages in the training phase: First, a warmup stage that learns to obtain a basic understanding of the geometric, semantic, and dynamic priors. Second, an initialization stage for the voxelized occupancy probabilities $\mathcal{V}^{\text{M}}$, involving the connected components algorithm. Third, an end-to-end optimization stage that jointly optimizes $\{\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}},\mathcal{V}^{\text{M}},\{\mathbf{z}_{n}\}_{n=1}^{N},f_{\psi}\}$ to refine the voxelized object occupancy.
Algorithm 1 Object-Centric Voxel Grids Initialization
1:  Input: $\{\sigma_{k}\}_{k=1}^{N_{s}}$, $\{\mathbf{x}_{k}\}_{k=1}^{N_{s}}$, $\{\mathbf{c}^{\text{ind}}_{k}\}_{k=1}^{N_{s}}$, $\{y_{k}\}_{k=1}^{N_{s}}$, hyperparameters $\delta_{\text{den}}$, $\delta_{\text{pos}}$, $\delta_{\text{rgb}}$, $\delta_{\text{vel}}$, $N$
2:  Output: $\mathcal{V}^{\text{M}}$
3:  $\mathcal{N}=\{k\mid\sigma_{k}>\delta_{\text{den}},k\in[1,N_{s}]\}$triangleright Filter out invalid locations
4:  for $t\xleftarrow{}1$ to $T-1$ do
5:     Calculate velocity $\mathbf{v}_{t,j}=f^{\prime}_{\xi}(\mathbf{x}_{j},t+1)-f^{\prime}_{\xi}(\mathbf{x}_{j},t)$ for $j\in\mathcal{N}$triangleright Get velocity information
6:  end for
7:  for Neighbor$(p,q)\in\{(u,v)\mid u,v\in\mathcal{N},l_{2}(\mathbf{x}_{u},\mathbf{x}_{v})<\delta_{\text{pos}}\}$ do
8:     $D_{\text{rgb}}(p,q)=l_{2}(\mathbf{c}^{\text{ind}}_{p},\mathbf{c}^{\text{ind}}_{q})$
9:     $D_{\text{vel}}(p,q)=\text{max}(l_{2}(\mathbf{v}_{t,p},\mathbf{v}_{t,q})$ for $t\xleftarrow{}1$ to $T-1)$
10:     if $D_{\text{rgb}}(p,q)<\delta_{\text{rgb}}$ and $D_{\text{vel}}(p,q)<\delta_{\text{vel}}$ and $y_{p}=y_{q}$  then
11:        $E(p,q)=1$triangleright Add $E(p,q)$ to the edge set $\mathcal{E}$
12:     else
13:        $E(p,q)=0$
14:     end if
15:  end for
16:  Generate the feature graph $G=(\mathcal{N},\mathcal{E})$
17:  Obtain the cluster set $\{\text{cl}_{n}\}_{n=1}^{M}=\text{ConnectedComponents}(G)$
18:  $\{\text{cl}_{n}\}_{n=1}^{N}=\text{sort-and-interp}(\{\text{cl}_{n}\}_{n=1}^{M})$triangleright Sort clusters and interpolate the rest
19:  for $n\xleftarrow{}1$ to $N$ do
20:     $\mathcal{V}^{\text{M}}[n,:,:,:]=\text{one-hot}(\text{cl}_{k})$triangleright Generate 4D voxel grids
21:  end for
22:  Rendering 2D masks $\widehat{P}(\mathbf{r})$ with $\mathcal{V}^{\text{M}}$
23:  Obtaining pseudo-labels by CRF: $\bar{P}(\mathbf{\mathbf{r}})=\text{CRF}(\widehat{P}(\mathbf{r}),C(\mathbf{r}))$
3.4.1 Warmup
To reduce the difficulty of learning the object-centric 4D voxel grids $\mathcal{V}^{\text{M}}$, we first use $\{\mathbf{I}_{t}\}_{t=1}^{T}$ and $\{\mathbf{F}_{t}\}_{t=1}^{T}$ to warm-up the density grids $\mathcal{V}^{\text{D}}$, the color-related feature grids $\mathcal{V}^{\text{C}}$, the semantic feature grids $\mathcal{V}^{\text{S}}$, the slot features $\mathbf{S}$, and the dynamics module $f_{\psi}(\cdot)$. Additionally, we train an extra forward dynamics module $f^{\prime}_{\xi}(\cdot)$ and a non-compositional neural renderer $N_{\phi^{\prime}}(\cdot)$. The overall objective is defined as $\mathcal{L}_{\text{Warm}}=\mathcal{L}_{\text{Render}}+\alpha_{p}\mathcal{L}_{\text{Point}}+\alpha_{b}\mathcal{L}_{\text{Bg-Ent}}+\alpha_{c}\mathcal{L}_{\text{Cyc}}+\mathcal{L}_{\text{Recon}}+\alpha_{e}\mathcal{L}_{\text{Ent}}$. This framework divides its focus between scene reconstruction, covered by the first four terms, and the representation of semantic features, addressed by the last two terms. The hyperparameter values of scene reconstruction are adopted from prior literature [22].
At a specific timestamp, we first use the rendering loss $\mathcal{L}_{\text{Render}}$ between the predicted and observed pixel colors, the background entropy loss $\mathcal{L}_{\text{Bg-Ent}}$, and the per-point RGB loss $\mathcal{L}_{\text{Point}}$ following DVGO [23] as basic objective terms for neural rendering:
$\begin{split}&\mathcal{L}_{\text{Render}}=\frac{1}{|\mathcal{R}|}\sum_{{r}\in\mathcal{R}}\left\|\widehat{C}(\mathbf{r})-C(\mathbf{r})\right\|_{2}^{2}\\ &\mathcal{L}_{\text{Bg-Ent}}=\frac{1}{|\mathcal{R}|}\sum_{{r}\in\mathcal{R}}-\widehat{w}_{l}^{r}\log(\widehat{w}_{l}^{r})-(1-\widehat{w}_{l}^{r})\log(1-\widehat{w}_{l}^{r})\\ &\mathcal{L}_{\text{Point}}=\frac{1}{|\mathcal{R}|}\sum_{{r}\in\mathcal{R}}\left(\frac{1}{P}\sum_{i=1}^{P}\left\|\overline{\mathbf{c}}_{i}-C(\mathbf{r})\right\|_{2}^{2}\right),\\ \end{split}$ (8)
where $\mathcal{R}$ is the set of sampled rays in a batch, $P$ is the number of sampling points along ray $r$, and $\widehat{w}_{l}^{r}$ is the color contribution of the last sampling point obtained by $\widehat{w}^{r}_{l}=T_{P}(1-\exp(-\sigma_{P}\delta_{P}))$. $\mathcal{L}_{\text{Bg-Ent}}$ can be viewed as a regularization to encourage the renderer to concentrate on either foreground or background. To enhance the coherence of the learned canonical-space dynamics, we train an additional module $f^{\prime}_{\xi}(\cdot)$ using the following cycle-consistency loss:
$\ \mathcal{L}_{\text{Cyc}}=\frac{1}{|\mathcal{R}|}\sum_{{r}\in\mathcal{R}}\left(\frac{1}{P}\sum_{i=1}^{P}\left\|f_{\psi}(x_{i},t)+f^{\prime}_{\xi}(x^{\prime}_{i},t)\right\|_{2}^{2}\right),$ (9)
where $x^{\prime}_{i}=x_{i}+f_{\psi}(x_{i},t),i\in[1,P]$. Furthermore, we incorporate the reconstruction loss $\mathcal{L}_{\text{Recon}}$ between the predicted and observed pixel semantic features, as well as the entropy loss $\mathcal{L}_{\text{Ent}}$ on the semantic probabilities to regularize them towards a one-hot distribution:
$\begin{split}&\mathcal{L}_{\text{Recon}}=\frac{1}{|\mathcal{R}|}\sum_{{r}\in\mathcal{R}}\left\|\widehat{F}(\mathbf{r})-F(\mathbf{r})\right\|_{1}\\ &\mathcal{L}_{\text{Ent}}=\frac{1}{|\mathcal{R}|}\sum_{{r}\in\mathcal{R}}\sum_{n=1}^{N}-\widehat{\mathcal{A}}(\mathbf{r})_{n}\log(\widehat{\mathcal{A}}(\mathbf{r})_{n}).\\ \end{split}$ (10)
3.4.2 Object-Centric Voxel Grids Initialization
We initialize the 4D object-centric grids $\mathcal{V}^{\text{M}}$ with two steps: (1) the feature graph generation and connected components computation and (2) a post-processing step with Conditional Random Field (CRF). As illustrated in Alg. 1, the entire process takes as inputs the voxel position set $\{\mathbf{x}_{k}\}$ with a size of $N_{s}=N_{x}\times N_{y}\times N_{z}$, the voxel density set $\{\sigma_{k}\}$ from $\mathcal{V}^{\text{D}}$, the view-independent colors $\{\mathbf{c}_{k}^{\text{ind}}\}$ generated by $N_{\phi^{\prime}}$, and the corresponding semantic labels $\{y_{k}\}$ obtained by $\mathrm{argmax}_{N}\left(\mathcal{A}\right)$ in Eq. (5). We also consider the forward canonical-space transitions generated by $f^{\prime}_{\xi}(\cdot)$. Besides, we set five hyperparameters in advance, where $\delta_{\text{den}}$ represents the density value threshold, $\delta_{\text{pos}}$ represents the position distance threshold, $\delta_{\text{rgb}}$ represents the RGB distance threshold, $\delta_{\text{vel}}$ represents the velocity distance threshold, and $N$ represents the maximum number of objects.
Specifically, we first filter out invalid locations from $\{\mathbf{x}_{k}\}$ with density values below the predefined threshold $\delta_{\text{den}}$. The remaining voxels are set as the nodes of a feature graph $G$. Next, we create the edges between adjacent graph nodes in $G$ based on their similarities in position, color, velocity, and semantics, where each node retains features of $(\mathbf{x}_{k},\mathbf{c}_{k}^{\text{ind}},\mathbf{v}_{1:T-1,k},y_{k})$ of the valid voxels. Subsequently, we perform the connected components algorithm on $G$, where the $M$ output clusters symbolize different objects, based on our assumption that voxels of the same object share similar motion, appearance, and semantics. If $M\leq N$, we directly utilize the clustering results and keep the remaining $N-M$ clusters empty. If $M>N$, we prioritize the clusters based on their contributions to the rendering results, as described in [24]. We initialize $\mathcal{V}^{M}$ by retaining the top $N$ clusters and applying nearest neighbor interpolation to the remaining voxel grids.
Furthermore, as an optional post-processing step in DynaVol-S, we employ a Conditional Random Field (CRF) to refine $\mathcal{V}^{M}$, following the common practice in video object segmentation, this step is beneficial for mitigating potential noise introduced by the connected components algorithm, thereby smoothing the segmentation results. The CRF method leverages RGB information to refine 2D mask predictions. Although this step could potentially compromise the 4D consistency of segmentation results, we address this by using the refined masks as pseudo-labels in the subsequent stage. This allows us to finetune the 4D voxel grids, $\mathcal{V}^{\text{M}}$, thereby ensuring consistency in object-centric representations.
3.4.3 Multi-Grids Joint Optimization
In this stage, we finetune the dynamics module $f_{\psi}$, the geometric grids $\mathcal{V}^{\text{D}}$, the appearance grids $\mathcal{V}^{\text{C}}$, and the object-centric grids $\mathcal{V}^{\text{M}}$ obtained in the first two stages. We finetune the parameters of $(\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}},\mathcal{V}^{\text{M}},\psi)$ and begin training ($(\{\mathbf{z}_{n}\}_{n=1}^{N},\phi)$ from scratch. Following an end-to-end training scheme, the dynamics module, object-centric latent codes, and compositional renderer collaboratively contribute to object-centric dynamic scene understanding. The loss function in this stage is defined as $\mathcal{L}_{\text{Dyn}}=\mathcal{L}_{\text{Render}}+\alpha_{p}\mathcal{L}_{\text{Point}}+\alpha_{b}\mathcal{L}_{\text{Bg-Ent}}$, complemented by a cross-entropy loss between the rendered 2D segmentations $\widehat{P}(\mathbf{r})$ from $\mathcal{V}^{\text{M}}$ and the pseudo-labels $\bar{P}(\mathbf{r})$ derived from the second stage: $\mathcal{L}_{\text{CE}}=-\sum_{n=1}^{N}\bar{P}(\mathbf{r})_{n}\log\widehat{P}(\mathbf{r})_{n}$.
4 Experiments
TABLE II: Novel view synthesis results of our approach compared with D-NeRF, DeVRF, TineuVox, and a variant of DeVRF aligned with our experimental configurations (see text for details). We show results averaged over $60$ novel views (one view per timestamp).
3ObjFall 3ObjRand 3ObjMetal 3Fall+3Still
[…]
D-NeRF 28.54 0.946 12.62 0.853 27.83 0.945 24.56 0.908
DeVRF 24.92 0.927 22.27 0.912 25.24 0.931 24.80 0.931
DeVRF-Dyn 18.81 0.799 18.43 0.799 17.24 0.769 17.78 0.765
TineuVox 32.79 0.968 29.46 0.955 30.84 0.960 29.61 0.950
DynaVol 31.83 0.967 31.10 0.966 29.28 0.954 27.83 0.942
DynaVol-S 32.15 0.963 32.58 0.973 30.87 0.963 31.59 0.966
6ObjFall 8ObjFall 3ObjRealSimp 3ObjRealCmpx
[…]
D-NeRF 28.27 0.940 27.44 0.923 27.04 0.927 20.73 0.864
DeVRF 24.83 0.905 24.87 0.915 24.81 0.922 21.77 0.891
DeVRF-Dyn 17.35 0.738 16.19 0.711 18.64 0.717 17.40 0.778
TineuVox 30.62 0.955 31.27 0.952 30.70 0.954 27.72 0.927
DynaVol 29.70 0.948 29.86 0.945 30.20 0.952 26.80 0.917
DynaVol-S 31.37 0.961 31.78 0.961 30.93 0.958 28.82 0.941
TABLE III: Novel view synthesis results using real-world scenes from HyperNeRF and $\text{D}^{2}$eRF. On average (as shown in the last column), our approach outperforms HyperNeRF by $10.8\%$ in PSNR and $8.3\%$ in SSIM. Additionally, DynaVol-S demonstrates better performance compared to $\text{D}^{2}$eRF, another neural renderer that incorporates disentanglement learning.
Chicken Broom Peel-Banana Duck Balloon Avg.
[…]
NeurDiff 21.17 0.822 17.75 0.468 19.43 0.748 21.92 0.862 20.13 0.836 20.08 0.747
HyperNeRF 26.90 0.948 19.30 0.591 22.10 0.780 20.64 0.830 17.81 0.803 21.35 0.790
$\text{D}^{2}$eRF 24.27 0.890 20.66 0.712 21.35 0.820 22.07 0.856 20.92 0.858 21.85 0.827
TineuVox 28.30 0.947 21.43 0.686 24.40 0.873 21.85 0.895 21.04 0.850 23.40 0.850
DynaVol 27.01 0.934 21.49 0.702 24.07 0.863 21.43 0.878 21.07 0.863 23.01 0.848
DynaVol-S 28.43 0.949 20.98 0.671 25.11 0.894 22.40 0.902 21.38 0.864 23.66 0.856
4.1 Experimental Setup
Implementation details. We set the size of voxel grids to $160^{3}$, the assumed number of maximum objects to $N=15$ for real-world scenes and $10$ for synthetic scenes, the dimension of semantic features to $D_{s}=16$, the dimension of latent codes to $D_{z}=64$, and the dimension of color-related features to $D_{c}=6$. We use $2$ hidden layers with $256$ channels in the renderer, and use the Adam optimizer with a batch of $4{,}096$ rays in the two training stages. The base learning rates are $0.1$ for the voxel grids and $1e^{-3}$ for all model parameters in the warmup stage and then adjusted to $0.01$ and $1e^{-4}$ in the second training stage. The two training stages both last for $20$k iterations. All experiments run on an NVIDIA RTX4090 GPU and last for about $2$ hours.
Datasets. We build the $8$ synthetic dynamic scenes in Table II using the Kubric simulator [25]. Each scene spans $60$ timestamps and contains different numbers of objects in various colors, shapes, and textures. The objects have diverse motion patterns and initial velocities. All images have dimensions of $512\times 512$ pixels. We also adopt real-world scenes from HyperNeRF [8] and $\text{D}^{2}$eRF [9], as shown in Table III and Table V. For the synthetic scenes, we follow D-NeRF [17] to employ images collected at viewpoints randomly sampled on the upper hemisphere. In contrast, real-world scenes are captured using a mobile monocular camera.
  • 3ObjFall. The scene consists of two cubes and a cylinder. Initially, these objects are positioned randomly within the scene and undergo a free-fall motion along the Z-axis.
  • 3ObjRand. We use random initial velocities along each object’s X and Y axes in 3ObjFall.
  • 3ObjMetal. We change the material of each object in 3ObjFall from “Rubber” to “Metal”.
  • 3Fall+3Still. We add another three static objects with complex geometry to 3ObjFall.
  • 6ObjFall & 8ObjFall. We increase the number of objects in 3ObjFall to $6$ and $8$.
  • 3ObjRealSimp. We modify 3ObjFall with real-world objects with simple textures.
  • 3ObjRealCmpx. We modify 3ObjFall with real-world objects that have complex textures.
  • Real-world data. We adopt the real-world scenes from HyperNeRF [8], $\text{D}^{2}$eRF [9], and NeRF-DS [26].
Metrics. For novel view synthesis, we report PSNR and SSIM [27]. To compare the scene decomposition results with the 2D methods, we employ the Foreground Adjusted Rand Index (FG-ARI) [28, 29]. It measures the similarity of clustering results based on the foreground object masks to the ground truth, ranging from $-1$ to $1$ (higher is better). Since no ground truth masks are provided in the real-world datasets, we employ the SAM-Tracker [30] to annotate one object in each scene. Additionally, we include the mean Intersection over Union (mIOU) to provide a comprehensive assessment of the segmentation accuracy and overlap with the ground truth. We utilize the first $5$ frames of the sequence to establish the correspondence between the predicted segmentation and the ground truth. This correspondence is then extrapolated across the entire sequence.
Figure 3: Novel view synthesis results. D-NeRF fails in the left synthetic examples on 3ObjRand. In the right real-world scenes, DynaVol-S significantly outperforms the prior art, TineuVox, by learning object-centric features.
TABLE IV: Scene decomposition results on synthetic datasets. To compare with SAVi, which takes as inputs video frames with a fixed viewpoint, we generate synthetic data with a stationary camera and evaluate DynaVol-S (FixCam) with fixed viewpoints.
Method 3D Decomp. 3ObjFall 3ObjRand 3ObjMetal 3Fall+3Still
[…]
SAVi (FixCam) 3.74 10.96 4.38 11.97 3.38 10.99 6.12 7.21
DynaVol-S (FixCam) 96.02 94.59 95.03 92.66 96.61 95.14 94.70 92.98
uORF 28.77 3.67 38.65 2.68 22.58 3.54 36.70 3.15
SAM 70.77 78.87 55.52 65.05 46.80 38.59 47.36 33.53
MovingParts 92.44 21.45 97.05 55.98 95.72 50.56 8.37 11.36
DynaVol 96.89 90.32 96.11 66.85 85.78 59.97 92.76 88.20
DynaVol-S 97.24 91.53 97.23 90.60 97.25 94.28 95.26 91.19
Method 3D Decomp. 6ObjFall 8ObjFall 3ObjRealSimp 3ObjRealCmpx
[…]
SAVi (FixCam) 6.85 16.65 7.87 19.44 3.10 14.10 4.82 9.98
DynaVol-S (FixCam) 94.40 93.04 94.72 86.23 95.54 87.75 94.34 92.88
uORF 29.23 3.87 31.93 2.53 38.26 2.85 33.76 3.39
SAM 62.66 41.35 71.68 28.52 56.65 57.46 51.91 75.20
MovingParts 78.06 11.79 89.83 22.25 92.34 36.77 87.07 19.91
DynaVol 95.61 89.59 93.38 87.16 94.28 64.75 95.02 63.56
DynaVol-S 96.08 89.80 95.60 85.41 95.37 84.55 96.55 92.61
Compared methods. In the novel view synthesis task for synthetic scenes, we evaluate DynaVol-S against D-NeRF [17], DeVRF [22], and TineuVox [7]. For real-world scenes, we compare it with NeuralDiff [14], HyperNeRF [8], TineuVox, and $\text{D}^{2}$eRF [9]. Furthermore, we specifically use TineuVox-Small for synthetic scenes and TineuVox-Base for real-world scenes to ensure alignment with our hyperparameter settings. In the scene decomposition task, we use established 2D/3D object-centric representation learning methods, SAVi [11], uORF [12], and MovingParts [13] as the baselines, where SAVi and uORF are pre-trained on the MOVi-A [25] and CLEVR-567 [12] datasets respectively, which are similar to our synthetic scenes. Besides, we also include a pretrained and publicly available Segment Anything Model (SAM) [31] as a competitive segmentation baseline. For real-world scenes, we compare DynaVol-S with OCLR [15] and ClipSeg [16].
TABLE V: Scene decomposition results on real-world datasets.
Chicken Torchocolate Keyboard Split-Cookie Americano Avg.
[…]
OCLR 14.31 14.80 22.46 20.21 6.54 10.30 34.39 34.41 46.15 35.32 24.77 23.01
OCLR (Test Adapt) 8.48 6.89 3.06 3.86 -5.51 1.63 17.20 14.02 5.78 3.59 5.80 5.99
ClipSeg 78.73 73.24 82.98 72.17 78.97 72.41 80.22 69.49 83.33 73.07 80.84 72.07
DynaVol 61.92 55.45 87.62 79.53 -5.52 0.42 -4.26 0.11 83.03 73.07 44.55 41.71
DynaVol-S 92.91 90.70 91.70 85.56 89.88 85.73 94.49 90.66 95.36 92.11 92.86 88.95
Scene decomposition. To get the 2D segmentation results, we assign the rays to different slots according to the contribution of each slot to the final color of the ray. Specifically, suppose $p_{in}$ is the occupancy probability of latent code $n$ at point $i$, we have
$\widehat{P}_{n}(\mathbf{r})=\sum_{i=1}^{P}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))p_{i,n},$ (11)
where $p_{in}$ is the corresponding occupancy probability and $\widehat{P}_{n}(\mathbf{r})$ is the color contribution to the final color of latent code $n$. We can then predict the label of 2D segmentation by $\hat{y}(\mathbf{r})=\mathrm{argmax}_{n}(\widehat{P}_{n}(\mathbf{r}))$.
Dynamic scene editing. The object-centric representations acquired through DynaVol-S demonstrate the capability for seamless integration into scene editing workflows, eliminating the need for additional training. By manipulating the 4D voxel grid $\mathcal{V}^{\text{D}}$ learned in DynaVol-S, objects can be replaced, removed, duplicated, or added according to specific requirements. For example, we can swap the color between objects by swapping the corresponding latent codes assigned to each object. Moreover, the deformation field of individual objects can be replaced with user-defined trajectories (e.g., rotations and translations), enabling precise animation of the objects in the scene.
4.2 Novel View Synthesis
Synthetic scenes. We compare the performance of DynaVol-S for novel view synthesis with D-NeRF, DeVRF, and TineuVox. As DeVRF is trained with $V=60$ views at the initial timestamp and $4$ views for subsequent timestamps, we also train a DeVRF model under a data configuration similar to ours. Specifically, we only have access to one image per timestamp ($t\geq 1$) in a dynamically sampled view from the upper hemisphere. This model is termed as “DeVRF-Dyn”. As shown in Table II, DeVRF-Dyn yields a significant decline in performance compared to its standard baseline, primarily due to its heavy reliance on accurate initial scene understanding. In contrast, DynaVol-S achieves the best results across most scenes among all the compared models.
Real-world scenes. We also evaluate the performance of DynaVol-S in real-world scenes with cutting-edge NeRF-based methods, including NeurDiff, HyperNeRF, TineuVox, and $\text{D}^{2}$eRF. As shown in Table III, DynaVol-S outperforms the previous state-of-the-art method, HyperNeRF, by $10.8\%$ in PSNR and $8.3\%$ in SSIM on average. Furthermore, we also compare DynaVol-S with another neural rendering method that learns disentangled representations. The results indicate that our approach outperforms $\text{D}^{2}$eRF by $8.2\%$ in PSNR and $3.5\%$ in SSIM.
Figure 4: Visualization of scene decomposition results for each object in the synthetic 6ObjFall scene.
Figure 5: Real-world scene decomposition results for Chicken and Torchocolates.
Figure 6: Real-world scene decomposition results for Cup and Sieve.
Qualitative results. Fig. 3 showcases the rendered images at an arbitrary timestamp from a novel view. On the synthetic dataset, it shows that DynaVol-S captures 3D geometries and the motion patterns of different objects more accurately than the compared methods. In contrast, D-NeRF fails to model the complex motion in 3ObjRand and generates blurry results in 3ObjRealCpmx. DeVRF, on the other hand, also fails to accurately model the trajectories of the moving object with complex textures, as highlighted by the red box in 3ObjRealCmpx. The real-world examples on the right side in Fig. 3 illustrate that DynaVol-S generates more high-quality novel view images than the prior art, TineuVox.
4.3 Scene Decomposition
Synthetic scenes. To obtain quantitative results for 2D segmentation as well as the scene decomposition maps, we assign the casted rays in volume rendering to different latent codes according to each latent code’s contribution to the ray’s final color. Table IV provides a comparison between DynaVol-S with 2D/3D unsupervised object-centric decomposition methods (SAVi, uORF, and MovingParts) and a pretrained SAM. For a fair comparison with SAVi, which works on 2D video inputs, we implement a DynaVol-S model using consecutive images with a fixed camera view (termed as FixCam). For uORF and SAM, as they are primarily designed for static scenes, we preprocess the dynamic sequence into $T$ individual static scenes as the inputs of these models and evaluate their average performance on the whole sequence. For MovingParts, which generates hierarchical clustering results by iteratively merging the slots with the most similar motions, we select the outcome that achieves the highest mIOU as its final result. The FG-ARI and mIOU results in Table IV show that DynaVol-S outperform all of the compared models by large margins.
Real-world scenes. We further evaluate the scene decomposition results of DynaVol-S in real-world scenes against the state-of-the-art unsupervised video segmentation method, OCLR [15]. Notably, OCLR utilizes DINOv1 features [32] for test-time adaptation, referred to as OCLR (Test Adapt), to enhance temporal consistency. Additionally, we compare DynaVol-S with the open-world object segmentation approach, ClipSeg [16]. Since ClipSeg relies on text prompts, we manually adjust the prompts to tune the model. It is important to highlight that ClipSeg employs the CLIP model [33] and is further trained with true segmentation in a supervised manner, whereas our approach is entirely unsupervised. As demonstrated in Table V, DynaVol-S consistently achieves the best decomposition results, significantly outperforming all compared methods.
Figure 7: Dynamic scene editing examples for real-world (top) and synthetic scenes (bottom). We optimize DynaVol-S on the original scene, manually edit its object-centric voxel grids (e.g., object removal and trajectory modification), and subsequently generate a new image sequence based on the edited voxel grids.
TABLE VI: Ablation studies of the key components and the training stages in DynaVol-S.
Method 3ObjRand 8ObjFall 3ObjRealCmpx
[…]
w/o forward deformation $f_{\xi}^{\prime}(\cdot)$ 32.15 90.49 31.55 75.91 29.05 92.78
w/o CRF 32.57 90.56 31.79 85.37 28.85 92.49
w/o Multi-grids joint optim. 30.48 90.24 29.03 84.46 26.81 92.37
Multi-grids joint optim. from scratch 28.86 14.73 29.01 5.82 25.64 11.09
Full model 32.58 90.60 31.78 85.41 28.82 92.61
Qualitative results. In Fig. 4, we randomly select a timestamp on 6ObjFall and present the object-centric decomposition maps of SAVi, uORF, SAM, MovingParts, and DynaVol-S. We have the following three observations from the visualized examples. First, compared with the 2D models (SAVi and SAM), DynaVol-S can effectively handle severe occlusions between objects in 3D space. It shows the ability to infer the complete shape of the objects, as illustrated in the $1$st and $5$th latent codes. Second, compared with the 3D decomposition uORF, DynaVol-S can better segment the dynamic scene by leveraging explicitly meaningful spatiotemporal representations, while uORF only learns latent representations for each object. Furthermore, DynaVol-S is shown to adaptively work with redundant latent codes, in the sense that the pre-defined number of latent codes can be larger than the actual number of objects in the scene. In such cases, the additional latent codes learn to disentangle noise in visual observations or learn to not contribute significantly to image rendering, enhancing the model’s flexibility and robustness. In Fig. 5, we showcase the real-world scene decomposition results of DynaVol-S compared with OCLR and ClipSeg. Our findings reveal that DynaVol-S produces clearer 3D segmentation for each object, especially at object boundaries, while also maintaining temporal consistency across the video frames. In contrast, OCLR struggles to preserve temporal consistency. DynaVol-S also outperforms its our previous work, DynaVol, indicating the importance and effectiveness of incorporating semantic features. More real-world examples can be found in Fig. 6 and in the Appendix Fig. 11– 13.
TABLE VII: The analysis of the impact of the number of latent codes ($N$) on the performance of our approach.
# Latent Codes 3ObjRand 8ObjFall 3ObjRealCmpx
[…]
5 32.53 90.69 31.73 42.80 28.82 92.60
10 (Final model) 32.58 90.60 31.78 85.41 28.82 92.61
15 32.59 90.67 31.79 85.46 28.86 92.59
TABLE VIII: Hyperparameter choice of the threshold value for splitting the foreground and background in object-centric voxel grids initialization stage, and the weight of per-point RGB loss ($\alpha_{p}$).
Hyperparameter 3ObjRand 8ObjFall 3ObjRealCmpx
[…]
$\delta_{\text{den}}=0.5$ 32.56 90.71 31.77 85.20 28.82 92.61
$\delta_{\text{den}}=2$ 32.57 90.65 31.74 88.41 28.83 92.61
$\alpha_{p}=0$ 30.00 88.66 30.64 68.31 28.16 92.45
$\alpha_{p}=0.1$ 30.95 31.61 31.15 23.89 28.89 84.25
$\delta_{\text{den}}=1,\alpha_{p}=0.01$ (Final model) 32.58 90.60 31.78 85.41 28.82 92.61
4.4 Dynamic Scene Editing
After the training period, the object-centric voxel representations learned by DynaVol-S can be readily used in downstream tasks such as scene editing without the need for additional model tuning. DynaVol-S allows for easy manipulation of the observed scene by directly modifying the object occupancy values within the voxel grids or switching the learned deformation function to a pre-defined one. This flexibility empowers users to make various scene edits and modifications. For instance, in the first example in Fig. 7, we remove the hand that is pinching the toys. In the second example, we modify the dynamics of the shoe from falling to rotating. More showcases are included in the Appendix.
4.5 Further Analyses
Ablation studies. We present ablation study results in Table VI. First, we can find that the forward deformation network $f_{\xi}^{\prime}(\cdot)$ is crucial to the final rendering and decomposition results. Furthermore, in the absence of the multi-grids joint optimization stage, the performance of DynaVol-S significantly degrades, highlighting the importance of refining the object-centric voxel representation with a latent-code-based renderer. Additionally, we conduct an ablation study that performs multi-grids joint optimization from scratch with randomly initialized $\mathcal{V}^{\text{M}}$ and $f_{\phi}(\cdot)$. The results clearly show that excluding the warmup stage substantially impacts the final performance, especially for the scene decomposition results.
The number of object-centric latent codes. First, we examine the impact of the number of latent codes ($N$), which indicates the maximum number of objects in the scene. According to Table VII, the model achieves comparable results as long as $N$ exceeds the number of objects in the scene. The presence of redundant latent codes has only a minor influence on both rendering and decomposition results. Additionally, we showcase the per-object rendering results in Appendix Fig. 14, demonstrating DynaVol-S’s robustness to variations in the number of latent codes. Our findings reveal that redundant latent codes typically remain unassigned, effectively addressing the potential issue of over-segmentation.
Figure 8: Visualization of semantic probabilities learned with different weights of the entropy loss ($\alpha_{e}$).
Hyperparameters for the loss functions. The hyperparameters in the loss functions are set to $\alpha_{p}=0.01$, $\alpha_{b}=0.01$, $\alpha_{e}=0.008$, $\alpha_{c}=1.0$. We study the hyperparameter choice of the threshold value for splitting the foreground and background in the warmup stage, and the weight ($\alpha_{p}$) of per-point RGB loss in Table VIII. It can be found that our model is robust to the threshold, and the performance of DynaVol-S remains largely unaffected with different thresholds. Furthermore, it can be found that $\alpha_{p}=0.01$ performs better than $\alpha_{p}=0.0$, indicating that a conservative use of $\mathcal{L}_{\text{point}}$ can improve the performance. A possible reason is that it eases the training process by moderately penalizing the discrepancy of nearby sampling points on the same ray. When the weight of $\mathcal{L}_{\text{point}}$ increases to $\alpha_{p}=0.1$, the performance of the method decreases significantly. Such a substantial emphasis on $\mathcal{L}_{\text{point}}$ is intuitively unreasonable and can potentially introduce bias to the neural rendering process, thereby negatively impacting the final results. For the entropy loss in Eq. (10), it encourages the semantic probabilities $\widehat{\mathcal{A}}$ learned in the warmup phase to approach a one-hot distribution. In Fig. 8, we visualize $\widehat{\mathcal{A}}$ learned with different values of $\alpha_{e}$. By increasing the weight of entropy loss $\alpha_{e}$, we prevent the model from over-segmenting the 3D scene with a redundant number of slot features. In other words, tuning $\alpha_{e}$ allows us to use a larger default number of $N$ while balancing between over-segmentation and under-segmentation.
Figure 9: Qualitative results of alternative designs for the slot attention module. From left to right: Visualization of reconstructed semantic features $\widehat{F}(\mathbf{r})$, corresponding semantic probabilities $\widehat{\mathcal{A}}(\mathbf{r})$, and scene decomposition results $\widehat{P}(\mathbf{r})$ for various design choices of the volume slot attention module. $F(\mathbf{r})$ denotes the 2D semantic features extracted by DINOv2.
TABLE IX: The analysis of using different semantic feature extractors.
Semantic Feature Extractor Chicken Torchocolate Keyboard Avg.
[…]
None (DynaVol) 61.92 55.45 87.62 79.53 -5.52 0.42 48.01 45.13
DINOv1 91.33 88.76 6.40 7.85 87.83 83.18 61.85 59.93
MaskClip 24.94 36.79 -1.64 0.28 -8.90 14.00 4.80 17.02
ViT 95.84 94.59 3.37 5.91 86.52 81.70 61.01 60.73
DINOv2 (in DynaVol-S) 92.91 90.70 91.70 85.56 89.88 85.73 91.49 87.33
Alternative designs for volume slot attention. In considering the design of the volume slot attention module, we perform the experiments shown in Fig. 9. A straightforward approach, referred to as “Baseline”, rewrites Eq. (5) by removing the slot attention operator, such that: $\mathbf{f}_{i}^{s}=\text{Interp}\left(\mathbf{x}_{i}+f_{\psi}(\mathbf{x}_{i},t),\mathcal{V}^{\text{S}}\right)$ and $\widehat{F}(\mathbf{r})=\sum_{i=1}^{P}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right)\mathbf{f}_{i}^{s}$. In the Baseline model, we otherwise obtain $\mathcal{A}_{i}$ by applying the connected components algorithm to the cosine similarities greater than $0.9$ between adjacent voxel features. We can find that the Baseline model tends to merge features into a single cluster in $\widehat{\mathcal{A}}$. As shown, the ambiguous estimation of $\widehat{\mathcal{A}}$ results in a noisy outcome for the learned semantic features, thereby undermining the final scene decomposition results presented by $\mathcal{V}^{M}$. Another alternative approach uses the Gumbel-softmax to replace the original softmax operation in Eq. (5), which ensures that each voxel is exclusively associated with a single slot, with $\mathcal{A}_{i}$ representing a one-hot vector. In Fig. 9, we visualize the generated semantic features $\widehat{F}(\mathbf{r})$, the corresponding semantic probabilities $\widehat{\mathcal{A}}(\mathbf{r})$ obtained from Eq. (6), and the scene decomposition results $\widehat{P}(\mathbf{r})$. As we can see, the Gumbel-softmax method leads to noisy decomposition results due to its hard assignment strategy. Our method, in contrast, effectively mitigates the overfitting of 2D semantic feature noise, resulting in clearer boundaries and more accurate decomposition outcomes. It is important to note that $\mathcal{A}_{i}$ plays a crucial role in initializing 4D voxel grids, $\mathcal{V}^{\text{M}}$, which in turn affects the final decomposition results. Specifically, if $\mathcal{A}_{i}$ merges features excessively, DynaVol-S will lose the insights provided by the semantic features, potentially degrading to DynaVol.
Alternative semantic feature extractors. In addition to DINOv2, our approach can be seamlessly integrated with other semantic feature extractors, such as MaskClip [34], ViT [35], and DINOv1 [32]. In Table IX, we explore the impact of these models. The results show that DINOv2 outperforms the other models. Aside from feature quality, these discrepancies may also stem from DINOv2’s strong capabilities in maintaining multi-view consistency [36].
5 Related Work
Unsupervised 2D scene decomposition. Most existing methods in this area [37, 38, 39, 40] use latent features to represent objects in 2D scenes. The slot attention method [21] extracts object-centric latents with an attention block and repeatedly refines them using GRUs [41]. SAVi [11] extends slot attention to dynamic scenes by updating slots at each frame and using optical flow as the training target. STEVE [6] improves SAVi by replacing its spatial broadcast decoder with an autoregressive Transformer. SAVi++ [5] improves SAVi by incorporating depth information, enabling the modeling of static scenes with camera motion. OCLR [15] adopts layered optical flow representations to identify multiple objects across the video. However, OCLR faces challenges in maintaining long-term temporal consistency when applied to real-world scenes.
Unsupervised 3D scene decomposition. Recent methods [4, 42, 43, 12, 44] combine object-centric representations with view-dependent scene modeling techniques like neural radiance fields (NeRFs) [45]. ObSuRF [43] adopts the spatial broadcast decoder and takes depth information as training supervision. uORF [12] extracts the background latent and foreground latents from an input static image to handle background and foreground objects separately. For dynamic scenes, Guan et al. [46] proposed to use a set of particle-based explicit representations in the NeRF-based inverse rendering framework, which is particularly designed for fluid physics modeling. Driess et al. [47] explored the combination of an object-centric auto-encoder and volume rendering for dynamic scenes, which is relevant to our work. However, different from our unsupervised learning approach, it requires pre-prepared 2D object segments. MovingParts [13] considers motion as an important cue for identifying parts and parameterizes the scene motion by tracking the trajectory of particles on objects under a Lagrangian view, facilitating part-based representation learning.
NeRF-based dynamic scene rendering. There is another line of work that models 3D dynamics using NeRF-based methods [17, 48, 22, 9, 49, 50]. D-NeRF [17] uses a deformation network to map the coordinates of the dynamic fields to the canonical space. Li et al. [48] extended the original MLP in NeRF to incorporate the dynamics information and determine the 3D correspondence of the sampling points at nearby time steps. Li et al. [50] achieved significant improvements on dynamic scene benchmarks by representing motion trajectories by finding 3D correspondences for sampling points in nearby views. DeVRF [22] models dynamic scenes with volume grid features [23] and voxel deformation fields. $\text{D}^{2}$eRF [9] presents a motion decoupling framework. Unlike our approach, it cannot segment multiple moving objects. NeuralDiff [14] extends the standard NeRF reconstruction of static scenes with two dynamic components: one for transient objects and another for actors in egocentric videos. HyperNeRF [8] extends the canonical space to a higher dimension to reconstruct topologically varying scenes with discontinuous deformations. TineuVox [7] proposes to use explicit voxel grids to model temporal information, accelerating the learning time for dynamic scenes. However, these methods do not achieve an object-centric representation for dynamic scenes.
6 Conclusion
In this paper, we presented DynaVol-S, an inverse graphics method designed to understand 3D dynamic scenes using object-centric volumetric representations. Our approach introduces two novel contributions: First, it introduces a set of object-centric voxel grids to represent 3D dynamic scenes, ensuring a view-consistent understanding of object geometries and physical interactions. Second, it integrates 2D semantic priors learned from large-scale image datasets into 3D rendering through the volume slot attention module, significantly enhancing the model’s ability to capture complex real-world scenes. DynaVol-S demonstrates remarkable performance compared to existing models in both unsupervised scene decomposition and novel view synthesis across synthetic and real-world scenarios. Additionally, our approach extends beyond existing neural rendering techniques for dynamic scenes by offering additional capabilities for dynamic scene editing.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62250062, 62106144), the Shanghai Municipal Science and Technology Major Project (Grant No. 2021SHZDZX0102), and the Fundamental Research Funds for the Central Universities.